Please help me solve this question and please give me the working out
If Tom was ten years younger he would be exactly one-fourth the age of Mark. If Tom was ten years older he would be exactly one-half the present age of Mark. How old is Tom?

Answer by lwsshak3(11628)

Please help me solve this question and please give me the working out
If Tom was ten years younger he would be exactly one-fourth the age of Mark. If Tom was ten years older he would be exactly one-half the present age of Mark. How old is Tom?
**
let t=tom's age
let m=mark's age
..
t-10=m/4
t+10=m/2
subtract
-20=m/4-m/2
-20=-m/4
m=80
t-10=m/4=80/4=20
t=30
Tom's age=30